# Multi-modal Generative AI: Multi-modal LLMs, Diffusions and the Unification

### v2

@@ III-C1 Text-to-Video Diffusion Models
The most intuitive approach to utilizing the text-to-image model is modifying the self-attention mechanism, which gets the text-to-video model without any additional parameters. Text2Video-Zero [103] is one of the pioneer works. Rather than random initial the latents of all frames independently, Text2Video-Zero only samples the latent code $z_{T}^{1}$ of the first frame and applies $\Delta t$ DDIM backward steps to obtain $z_{T^{\prime}}^{1}$. After that, Text2Video-Zero determines the global scene and a camera motion direction, proposes a warping function $W_{k}$ to get all $F$ frames from $z^{1}_{T^{\prime}}$ to $z^{F}_{T^{\prime}}$, and then performs a DDPM forward to get the initial latents. To keep the consistency among different frames, Text2Video-Zero proposes cross-frame attention which uses keys and values from the first frame to generate the images. Latent-Shift [104] is another representative method. It proposes a novel Temporal-Shift module that splits the latents along the channel dimension and shifts the split channel along the temporal dimension to keep the consistency of all frames. These methods have fully used the powerful pretrained text-to-image models and can generate videos with much higher resolution and quality than traditional text-to-video methods using GANs and VAEs. However, rather than capturing, training, and understanding the temporal information, these methods are more like to provide a class of expert knowledge that can utilize the temporal information from a human perspective. Thus, these methods enjoy high generation efficiency but the videos generated still struggle with motion smoothness, dynamic degree, and video consistency.
@@
To solve the problems, another kind of approaches [105, 106, 107] not only inherits the architecture of the T2I models but also makes efforts to introduce novel modules or modify the original structure to learn the temporal information. VDM [105] is one of the earliest works that transferred the T2I model to solve T2V tasks. VDM proposes a 3D U-Net that modifies the diffusion architecture by changing each 2D spatial convolutional layer into a 3D convolution. After that, for each spatial attention block, VDM inserts a temporal attention block that performs attention over all frames with relative position embeddings to distinguish the ordering of frames. Make-a-video [106] proposed a pseudo-3D convolutional and attention layer which consists of a spatial 2D convolutional layer and a temporal 1D convolutional layer. Compared to 3D convolution, this approach is much more efficient while facilitating information sharing between the spatial and temporal axes. To more flexibly apply the capabilities of the T2I model such as the customization and style transferring ability brought by LoRA, AnimateDiff [107] keeps the original architecture and only inserts a motion module after each pretrained layer. The motion module consists of an input projection layer, several temporal self-attention layers, and an output projection layer. To avoid harming the original capabilities of T2I models, AnimateDiff zero initializes the output projection layer.
@@
At this point, the basic text-to-video models have been constructed based on the text-to-image model. However, there are still two problems. The first problem lies in the fact that these methods can only control the generation of video through text, and it is usually difficult to describe all aspects of the requirements of the video in text. How to better control the generation of video is an important issue. The second problem lies in the fact that, limited by the scale of model parameters and GPU memory, most of the videos generated by these methods are in the range of 16-24 frames, which makes it difficult to satisfy the needs of real-life users of needs of visual content. Next, we will analyze these two issues and discuss some related works.
III-C2 Controllable Generation with Diffusion Model
For controllable generation, the key challenge is how to choose suitable conditioning information and how to utilize this information fully.
An intuitive way to do this is to use existing videos, which can be regarded as video editing [110, 111, 112, 113]. StableVideo [111] first introduces a pretrained model to split foreground and background and edits them separately. To better maintain the consistency of subjects, StableVideo proposes an inter-frame propagation mechanism that utilizes the layered representations to keep information between different frames. Rerender-A-Video [112] proposes a novel two-stage editing approach. In the first stage, Rerender-A-Video identifies the keyframes of the reference video and edits them according to the given prompt. To ensure effectiveness, Rerender-A-Video introduces a pretrained image diffusion model with hierarchical cross-frame constraints. The second stage utilizes the edited keyframes to perform overall video editing through temporal-aware patch matching and frame blending. FateZero [113] makes full use of the information provided by the attention maps during inversion. On the one hand, they encapsulate a wealth of layout and motion information from the original video. On the other hand, through the cross-attention maps, a novel blending mask can be derived. These masks indicate the information that influences the subject requiring editing, thereby minimizing semantic loss during subject editing.
It can be found that these methods usually utilize two aspects of the reference video. One is the overall layout information of the video, including the position of the object, the motion trajectory, etc. The other is the attributes of the subjects requiring editing, with appropriate extraction and adjustment. This also implies that not all information in the video is valid according to the situation, such as the color or shape of an object, which often proves to be disruptive when we intend to edit it. With such considerations, certain approaches initially involve pre-extracting conditional information through auxiliary networks, and then feeding this preprocessed information into the generative model, aiming for improved controllability over video generation.
Inspired by controllable generation methods used in text-to-image tasks, ControlNet [114] is introduced to text-to-video generation [115, 116, 117]. It utilizes the information extracted from each frame of videos, such as skeleton, depth map, and optical flow, to generate videos that satisfy the provided text prompt. Control-A-Video [115] utilizes a ControlNet to control the generation process by different types of conditional information. Besides, to improve the consistency, Control-A-Video proposes a novel residual-based noise initialization strategy that introduces motion prior to the diffusion process. VideoControlNet [116] proposes a motion-guided method that uses an auxiliary model to predict the optical flow between keyframes. After generating the keyframes, VideoControlNet utilizes a motion-guided B-frame interpolation module to generate the rest frames. In contrast to the aforementioned methods, SparseCtrl [117] takes into account the potential quality degradation that arises from using noisy latents as inputs in traditional ControlNet. Therefore, SparseCtrl proposes a novel sparse condition encoder that proposes a sparse mask, eliminating the noisy sample input and exclusively accepting condition information. Other methods [118, 119] analogize condition information into embeddings and employ attention mechanisms to achieve controllable generation. Follow-Your-Pose [118] proposes a two-stage training strategy. In the first stage, Follow-Your-Pose trains a pose encoder to translate the frames of key points into specific embeddings. The second stage introduces a temporal self-attention and a cross-frame self-attention to keep consistency. During inference, Follow-Your-Pose mixes the embeddings of poses and the latents to control the video through the key points.
However, whether directly utilizing videos as control conditions or extracting crucial information from them, the aforementioned approaches heavily rely on the inherent structure of the reference videos being highly consistent with the desired generated videos. This limitation constrains the diversity in controllable video generation unless we possess an infinite video library and exceptionally powerful video retrieval methods to cater to all the whimsical imaginings of users. Hence, other works also explore the use of simpler and more readily accessible control conditions to exert finer control over specific aspects of videos.
Some methods attempt to control videos by emulating how people shoot movies in real life, such as controlling the layouts [120, 121], adjusting camera views [107, 122], and setting “actors” [123, 124]. Compared to previous control methods, the major advantage of these approaches lies in the simplicity of their condition information, such as sequences of bounding boxes, a viewpoint with a direction, or images of objects. Simultaneously, they can effectively represent a specific set of significant attributes of a video. This eliminates the need to invest substantial efforts in searching for a suitable reference video before generating the video. Users can obtain control information through their own understanding, either via a GUI for dragging and dropping or in numerous method. This not only ensures controllability but also significantly enhances the diversity of video generation. Next, we will briefly introduce several works of these methods.
LVD [125] analyzes potential layout information within the user-provided prompt through LLM and transforms it to frame-level dynamic scene layout (DSL) information, which can be seen as a series of bounding boxes. To use this DSL information for generating videos that satisfy the desired layouts, LVD introduces a DSL-guided video generator. It designs an energy function to assess the degree of overlap between the generated objects and the required bounding boxes and influences the object positions during denoising by minimizing the energy function and back-propagation. CameraCtrl [126] proposes a novel plug-and-play module that can control the camera trajectory for text-to-video generation. It trains a camera encoder that can output multi-scale camera representations, which are then utilized by the temporal attention layers of U-Net to control the video generation process. DisenStudio [124] addresses the challenge of customized multi-subject text-to-video generation in the real world where each subject has only a few reference images available. It proposes a disentangled spatial control approach to associate each subject with the desired action. Besides, DisenStudio proposes a novel multi-subject co-occurrence tuning and masked single-subject tuning to keep the visual attributes of given subjects, and a multi-subject motion-preserved tuning method to maintain the temporal motion-generation ability. Kaleido [127] integrates this condition information by encoding different control conditions to tokens, enabling more flexible multi-condition controllable generation. However, employing control conditions beyond text inevitably leads to potential conflicts among multiple conditions, resulting in a decline in the quality of the generated videos. This problem could serve as a challenging future direction.
III-C3 Long Video Generation
Another challenge of the diffusion model is generating longer videos. Some approaches [115, 117, 128, 129] leverage the controllable generation methods mentioned before, splitting the whole video into several smaller video chunks and generating them in an autoregressive manner. Typically, they use the final frames of the preceding chunk as a reference and generate the next chunk to ensure complete the same in the overlapping parts between chunks, thereby guaranteeing consistency and smoothness across different chunks. Other than using additional modules to fully control the overlapping frames between chunks, FreeNoise [130] accomplishes the generation of long videos by performing specific operations on the latents of overlapping frames between different chunks. FreeNoise no longer initializes noise for all frames but rearranges the noise sequences to achieve long-range correlations, and applies temporal attention through window-based fusion. The limitation of this kind of method is that it often suffers from concept drift as the video becomes longer. Additionally, it often fails to generate new backgrounds or videos with high dynamic degrees.
Another way to generate longer videos is to rely on larger datasets and model parameters. Early works pretrain U-Net based diffusion models [131, 132] which can only generate 1-second or 2-second videos. More recently, by scaling the DiT architecture for video generation, several works [2, 133, 134, 135] can generate videos up to 1 minute, with high resolution and smoothness. Despite their success, the length of the generated video cannot be arbitrarily long because of the computational restrictions. A possible way to generate longer videos is to train a multi-modal video generation model, which can receive the last frame or the last several frames of the previous video as input, and also the text prompt as input, to generate the next video clip.
@@ IV Unification of Understanding and Generation
Till now, we have discussed both the multi-modal LLMs and the multi-modal diffusion models, where the former works well for multi-modal understanding and the latter exhibits powerful ability in visual generation. Then there arises a natural question, could we have a unified model that can simultaneously work well for multi-modal understanding and generation? Next, we will discuss this trending problem from the following two perspectives: (i) the probabilistic modeling method, and (ii) the model architecture.
@@ IV-A1 Autoregressive (AR) Model
Although diffusion models have become dominant in visual generation, there are still some recent attempts [136, 3, 137, 138, 44, 139] on generating visual content in an autoregressive manner. These works will first try to map the input images and text into discrete tokens respectively. Particularly, the images are discretized with visual tokenizers such as VQGAN or VQ-VAE. Then the mixed text and visual tokens will be fed into a multi-modal autoregressive model. After that, the model will output the mixed text and visual tokens. Also, some special tokens such as $<soi>,<eoi>$ are used to indicate the start of the image tokens and the end of the image tokens. Then the generated text tokens will deliver how the model understands the input multi-modal information, and the visual tokens will be sent to the decoder of the VQ-VAE or VQGAN to reconstruct images. Therefore, the autoregressive model can be used for both understanding and visual generation.
Remark. Despite these efforts, the autoregressive method is far from perfect — it basically assumes the existence of a causal structure and causal attention, where previous tokens are used to predict next tokens. However, this is not suitable for image generation because it is difficult to determine which visual token should be the first and which one should be the last. Therefore, a recent work [140] tries to use the next-scale prediction paradigm to generate images, where the lower-resolution images are regarded as previous tokens to predict (next) higher-resolution images. Unfortunately, the scaling ability is still not verified in the multi-modal understanding and generation, and the model achieves 1.73 FID score on the ImageNet [141] benchmark for generation, falling behind the diffusion model [142] with 1.35 FID score.
@@ IV-A2 Joint Autoregressive and Diffusion Model
The first one is that we have a pretrained diffusion model for visual generation and a multi-modal LLM for multi-modal understanding. Then we connect these two parts and can obtain a unified model. About how to connect these two parts, many existing works [143, 144, 145] directly use the LLM as the controller, and the diffusion model as a tool for visual generation, which is a common paradigm in tool learning. Although works like tool learning can enable the models with visual generation abilities, they easily suffer generation failure when meeting multi-modal generation conditions. For example, when we want to generate “a specific girl (described with a given image) and a specific dog (described with a given image) playing on the grass”, the tools available are only SOTA text-to-image models. They will fail to guarantee the specific girl and dog occur in the generated image. In fact, there are many conditions that can not be described with only text, and this kind of tool-learning method will fail. To tackle the problem, a more advanced way is to train a learnable connector [146, 147, 148, 149, 150], which aligns the diffusion model and the multi-modal LLM in the same space, similar to the training paradigm of the alignment module in multi-modal LLM. The alignment process enables the diffusion model to receive the LLM output multi-modal embeddings as conditions instead of pure text descriptions, thus achieving multi-modal generation. However, this paradigm inherits the limitations of alignment architecture. The multi-modal LLM and the diffusion model are pretrained respectively, the performance of the unified model will be limited by each model. Additionally, from an intuitive perspective, multi-modal understanding and multi-modal generation should not be independent tasks, but two related tasks that could share knowledge. For example, when generating a picture of a girl riding a horse, the model should definitely understand the concepts “girl”, “horse” and “riding”. Therefore, although it is resource-friendly to obtain a unified multi-modal understanding and generation framework through connecting pretrained models, its ability is limited by independent modeling.
@@
The second possible model is a unified multi-modal-transformer framework as shown in Fig. 9, where we do not rely on two pretrained models, but try to use a single model trained with both diffusion and autoregressive regularizations. The multi-modal input processor will first transform the multi-modal data into sequences that can be received by the transformers. Then the multi-modal transformer will try to learn the multi-modal knowledge for both understanding and generation. The diffusion regularization is used to guide visual generation and the autoregressive regularization is used to guide the text generation. Note that this is a transformer-like model but not necessarily an LLM. This is because when using transformers to generate visual content, the full-attention mechanism is usually adopted. In contrast, the attention mechanism adopted by LLM is causal and uni-directional. Therefore, an adaptive or mixed attention mechanism might be designed. This perspective is verified in TransFusion [151] and Show-o [152]. The difference between Transfusion and Show-o mainly lies in the diffusion model, where TransFusion adopted continuous diffusion that is similar to current visual diffusion models, but Show-o adopted masked generative modeling [153] which could be regarded as discrete diffusion regularization. Therefore, Show-o still relies on a pixel-level visual tokenizer for image generation but might trade off some understanding ability. Additionally, these two works are primary attempts at combining autoregressive and diffusion modeling methods in a single transformer-like model. There still exist several open problems regarding what the model architecture should be like, such as the multi-modal input processor or the transformer-like model, which we will discuss next.

### v3

@@ III-C Text-to-Video Generation
The most intuitive approach to utilizing the text-to-image model is modifying the self-attention mechanism, which gets the text-to-video model without any additional parameters. Text2Video-Zero [99] is one of the pioneer works. Rather than randomly initializing the latents of all frames independently, Text2Video-Zero only samples the latent code $z_{T}^{1}$ of the first frame and applies $\Delta t$ DDIM backward steps to obtain $z_{T^{\prime}}^{1}$. After that, Text2Video-Zero determines the global scene and a camera motion direction, proposes a warping function $W_{k}$ to get all $F$ frames from $z^{1}_{T^{\prime}}$ to $z^{F}_{T^{\prime}}$, and then performs a DDPM forward to get the initial latents. To keep the consistency among different frames, Text2Video-Zero proposes cross-frame attention, which uses keys and values from the first frame to generate the images. Latent-Shift [100] is another representative method. It proposes a novel Temporal-Shift module that splits the latents along the channel dimension and shifts the split channel along the temporal dimension to keep the consistency of all frames. These methods have fully used the powerful pretrained text-to-image models and can generate videos with much higher resolution and quality than traditional text-to-video methods using GANs and VAEs. However, rather than capturing, training, and understanding the temporal information, these methods are more like providing a class of expert knowledge that can utilize the temporal information from a human perspective. Thus, these methods enjoy high generation efficiency, but the videos generated still struggle with motion smoothness and video consistency.
@@
To solve the problems, another kind of approaches [101, 102, 103] not only inherits the architecture of the T2I models but also makes efforts to introduce novel modules or modify the original structure to learn the temporal information. VDM [101] is one of the earliest works that transferred the T2I model to solve T2V tasks. VDM proposes a 3D U-Net that modifies the diffusion architecture by changing each 2D spatial convolutional layer into a 3D convolution. After that, for each spatial attention block, VDM inserts a temporal attention block that performs attention over all frames with relative position embeddings to distinguish the ordering of frames. Make-a-video [102] proposed a pseudo-3D convolutional and attention layer, which consists of a spatial 2D convolutional layer and a temporal 1D convolutional layer. Compared to 3D convolution, this approach is much more efficient while facilitating information sharing between the spatial and temporal axes. To more flexibly apply the capabilities of the T2I model, such as the customization and style transferring ability brought by LoRA, AnimateDiff [103] keeps the original architecture and only inserts a motion module after each pretrained layer. The motion module consists of an input projection layer, several temporal self-attention layers, and an output projection layer. To avoid harming the original capabilities of T2I models, AnimateDiff zero initializes the output projection layer.
@@
III-D Text-to-Speech Generation
Text-to-Speech (TTS) generation, also known as speech synthesis, is one of the most fundamental tasks in multimodal speech processing [106]. The development of TTS has evolved from a three-stage pipeline to a two-stage framework, and more recently, to end-to-end systems. Before the advent of neural networks, TTS systems typically converted text into linguistic features and then into acoustic features before decoding them into waveforms. With the introduction of neural networks, this process was simplified, where text only needs to be transformed into either linguistic or acoustic representations. Most recent diffusion-based TTS models adopt a two-stage approach: an acoustic model first generates acoustic features, which are then converted into waveforms using a vocoder. Moreover, several studies explore end-to-end TTS frameworks that directly synthesize speech waveforms from text input.
For two-stage text-to-speech diffusion models, the acoustic model and vocoder are the two key components. The acoustic model converts text into acoustic representations, while the vocoder synthesizes waveforms from these features. DiffWave [107] is one of the earliest diffusion-based speech synthesis models, serving as a neural vocoder. It formulates waveform generation as a DDPM task, where a neural network learns to reverse a gradual noising process applied to real waveforms. WaveGrad [108] also functions as a vocoder, introducing a continuous-time, score-based diffusion approach that models a gradient field to guide the denoising process, rather than relying on a discrete noise schedule. Grad-TTS [109] is a diffusion-based acoustic model that extends diffusion modeling from vocoders to full TTS systems. It generates acoustic features from text through stochastic differential equations (SDEs), enabling a non-autoregressive acoustic modeling framework. Diff-TTS [110] is another diffusion-based acoustic model that further advances speech synthesis by formulating the entire acoustic modeling process as a deterministic or stochastic denoising procedure.
Compared with two-stage approaches, end-to-end text-to-speech diffusion models reduce error propagation and produce higher-quality speech, becoming the mainstream development direction. For example, WaveGrad 2 [111] discards the two-stage design of WaveGrad [108] and adopts an end-to-end framework that directly generates audio from a phoneme sequence. Moreover, recent systems such as TTS-1 [112] and MiniMax-Speech [113] also follow end-to-end architectures and achieve remarkable performance in speech generation.
@@ IV Unification of Understanding and Generation
Until now, we have discussed both the multi-modal LLMs and the multi-modal diffusion models, where the former works well for multi-modal understanding and the latter exhibits a powerful ability in visual generation. Then a natural question arises: could we have a unified model that can simultaneously work well for multi-modal understanding and generation? Next, we will discuss this trending problem from the following two perspectives: (i) the probabilistic modeling method, and (ii) the model architecture.
@@ IV-A1 Autoregressive (AR) Model
Although diffusion models have become dominant in visual generation, there are still some recent attempts [114, 3, 115, 116, 117, 47, 118] on generating visual content in an autoregressive manner. These works will first try to map the input images and text into discrete tokens, respectively. Particularly, the images are discretized with visual tokenizers such as VQGAN or VQ-VAE. Then the mixed text and visual tokens will be fed into a multi-modal autoregressive model. After that, the model will output the mixed text and visual tokens. Also, some special tokens such as <soi>,<eoi><soi>,<eoi> are used to indicate the start of the image tokens and the end of the image tokens. Then the generated text tokens will deliver how the model understands the input multi-modal information, and the visual tokens will be sent to the decoder of the VQ-VAE or VQGAN to reconstruct images. Therefore, the autoregressive model can be used for both understanding and visual generation.
Remark. Despite these efforts, the autoregressive method is far from perfect — it basically assumes the existence of a causal structure and causal attention, where previous tokens are used to predict next tokens. However, this is not suitable for image generation because it is difficult to determine, which visual token should be the first and which one should be the last. Therefore, a recent work VAR [119] tries to use the next-scale prediction paradigm to generate images, where the lower-resolution images are regarded as previous tokens to predict (next) higher-resolution images. Unfortunately, the scaling ability is still not verified in multi-modal understanding and generation, and the model achieves a 1.73 FID score on the ImageNet [120] benchmark for generation, falling behind the diffusion model [121] which has a 1.35 FID score. In general, joint AR and diffusion models outperform unified AR models on visual generation tasks. For instance, EMU3 [47] and Janus-Pro [122], both unified AR models, achieve 0.66 and 0.80 on the GenEval benchmark, respectively. In contrast, joint AR-diffusion models such as Mogao [123] and Bagel [124] reach 0.89 and 0.88, demonstrating the advantages of combining AR and diffusion components for visual generation.
@@ IV-A2 Joint Autoregressive and Diffusion Model
The first one is that we have a pretrained diffusion model for visual generation and a multi-modal LLM for multi-modal understanding. We then connect these two components, forming what we call Connector-based Joint Models. Regarding how to connect these two parts, many existing works [125, 126, 127] directly use the LLM as the controller and the diffusion model as a tool for visual generation, which is a common paradigm in tool learning. Although works like tool learning can enable the models with visual generation abilities, they easily suffer from generation failure when meeting multi-modal generation conditions. For example, when we want to generate “a specific girl (described with a given image) and a specific dog (described with a given image) playing on the grass”, the tools available are only SOTA text-to-image models. They will fail to guarantee that the specific girl and dog occur in the generated image. In fact, there are many conditions that cannot be described with only text, and this kind of tool-learning method will fail. To tackle the problem, a more advanced way is to train a learnable connector [128, 129, 130, 131], which aligns the diffusion model and the multi-modal LLM in the same space, similar to the training paradigm of the alignment module in multi-modal LLM. The alignment process enables the diffusion model to receive the LLM output multi-modal embeddings as conditions instead of pure text descriptions, thus achieving multi-modal generation. However, this paradigm inherits the limitations of alignment architecture. The multi-modal LLM and the diffusion model are pretrained respectively. The performance of the unified model will be limited by each model. Additionally, from an intuitive perspective, multi-modal understanding and multi-modal generation should not be independent tasks but rather two related tasks that could share knowledge. To train such a model, both the MLLM and the diffusion model can be frozen, and only the connector is trained. This maximally preserves the capabilities of the two models, but the information bottleneck between them can be particularly severe. Alternatively, one or both of the models can be included in training, but this requires a larger amount of data and computational resources to ensure that the original abilities of the models are not compromised. For example, in Qwen-Image [132], the MLLM is kept frozen while the diffusion model is trained on a large dataset. This preserves the full capability of the MLLM while endowing it with strong generative ability.
@@
The second possible model is a unified multi-modal-transformer framework as shown in Fig. 8, where we do not rely on two pretrained models, but try to use a single model trained with both diffusion and autoregressive regularizations, which we refer to as Autoregressive-Diffusion Joint Models. The multi-modal input processor will first transform the multi-modal data into sequences that can be received by the transformers. Then the multi-modal transformer will try to learn the multi-modal knowledge for both understanding and generation. Specifically, the training objectives are designed differently for each modality: text prediction uses an autoregressive regularization (computed token-wise), while image prediction uses a diffusion regularization (computed over the entire image, covering multiple patches). During inference, the model dynamically switches between language modeling and diffusion modes. In language modeling mode, it samples tokens sequentially; upon generating the BOI token, it switches to diffusion mode, appending a sequence of pure noise patches corresponding to the target image size, and gradually generates the image through T-step denoising iterations. At each step, the model predicts the noise based on the current image representation and updates the patch sequence until denoising is complete. The EOI token is then appended, and the model switches back to language modeling mode. Note that this is a transformer-like model but not necessarily an LLM. This is because when using transformers to generate visual content, the full-attention mechanism is usually adopted. In contrast, the attention mechanism adopted by LLM is causal and uni-directional. Therefore, an adaptive or mixed attention mechanism might be designed. This perspective is verified in TransFusion [133] and Show-o [134]. The difference between Transfusion and Show-o mainly lies in the diffusion model, where TransFusion adopts continuous diffusion that is similar to current visual diffusion models, but Show-o adopts masked generative modeling [135], which could be regarded as discrete diffusion regularization. Therefore, Show-o still relies on a pixel-level visual tokenizer for image generation but might trade off some understanding ability. Additionally, these two works are primary attempts at combining autoregressive and diffusion modeling methods in a single transformer-like model. There still exist several open problems regarding what the model architecture should be like, such as the multi-modal input processor or the transformer-like model, which we will discuss next.